OpenStudy (lilkg77):

x2 + 4x = 14 A. (x + 4)2 = 30 B. (x + 4)2 = 18 C. (x + 2)2 = 18 D. (x + 2)2 = 30

OpenStudy (lilkg77):

x2 is a power

OpenStudy (gabbyalicorn):

Equation: x2+4x=14 Subtract 14 from both sides: x2+4x−14=14−14 x2+4x−14=0 Use quadratic formula with a=1, b=4, c=-14: x=−b±b2−4ac2a x=−(4)±(4)2−4(1)(−14)2(1) x=−4±722 x=−2+32 or x=−2−32 Hopefully I did this right. Answer: x=−2+32 or x=−2−32

OpenStudy (lilkg77):

This cant be right ^

OpenStudy (lilkg77):